6) A company ships computer components in boxes that contain 10 items. Assume that the6) probability of a defective computer component is0.1. Find the probability that in a sample of 5 items the first defective item is the fourth or fifth one checked.
Answer #1

Let X denotes the number of trials that are needed to find the first detective item.

The probability that first defective item is the fourth or fifth one checked

= P(X = 4) + P(X = 5)

= [(1-0.1)^3*0.1] + [(1-0.1)^4*0.1]

(0.9 *0.1(0.94 0.1)

0.0729 0.06561

-0.138510